Production priorties

Under:


(Weekly update from production team)

1. Data production
(1) Run18 Isobar final production 
- done
(2) BES-II: Run19 19.6 COL > Run19 14.6 COL > Run20 31.2 FXT (7.7) - pending TPC misalignment calibration
(3) Run17 pp510 st_W > st_physics - st_W done, st_physics ongoing
(4) Run19 BES-II: 4.59 FXT>200 COL > 31.2 FXT (7.7)
(5) Run17 pp510 st_dimuon
 using SL20a or beyond, requested by HF- vertex selection based on MTD matching developed. Will need new library before start.
(5) Run17 pp510 st_rp to apply latest space charge correction - ongoing
(6) Run20 BES-II: 19.5 FXT (6.2) > 13.5 FXT (5.2) > 9.8 FXT (4.5) > (Run19+20) 7.3 FXT (3.9) > 5.75 FXT (3.5) > 9.2 COL > 11.5 COL > 26.5 FXT (7.2)
(7) Run05 Cu+Cueast-west asymmetry in produced data


2. MuDst to picoDst conversion

(1) Run14 AuAu 200 GeV st_physicsP18ih-MuDst using SL20a or beyond, requested by JetCorr - done
mtdMatch, y2014a, picoDst,PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:3, PicoCovMtxMode:PicoCovMtxWrite, PicoBEmcSmdMode:PicoBEmcSmdWrite
(2) pp200_production_2012, P12id-MuDst using SL20b or beyond, requested by JetCorr
PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:6, PicoCovMtxMode:PicoCovMtxWrite, PicoBEmcSmdMode:PicoBEmcSmdWrite
(3) 2010 AuAu200_production, UU_production_2012:
picoDst PicoVtxMode:PicoVtxDefault PicoCovMtxMode:PicoCovMtxWrite PicoBEmcSmdMode:PicoBEmcSmdWrite
(4) BES-I data:
Tentative chain opinion:
 picoDst PicoVtxMode:PicoVtxDefault PicoCovMtxMode:PicoCovMtxWrite PicoBEmcSmdMode:PicoBEmcSmdWrite
(5) Run14 AuAu 200 GeV st_mtd, requested by HF
mtdMatch, y2014a, PicoVtxMode:PicoVtxVpdOrDefault, TpcVpdVzDiffCut:3, PicoCovMtxMode:PicoCovMtxSkip


New request:


3. PicoDst restoration (MuDst is specified) (Progress report)
(1) Isobar production, P20ic_SL20c, st_physics
 - done
(2) pp500_production_2013, P14ia, st_physics and st_physics_adc, MuDst - done
- Run list: /star/u/agibson/2013Pi0s/2020-December-Run13PeriodWishList.txt
(3) cuProductionMinBias P17ii SL18b 
- done
(4) AuAu_200_production_2016, st_physics, P16ij_SL20c 
(5) AuAu200_production2_2016, st_sst, st_nosst, P16ij_SL19c 
(6) production_pp200long_2015, production_pp200long2_2015, production_pp200trans_2015, st_mtd, P16id_SL19c
(7) AuAu_200_production_2014, AuAu_200_production_low/mid/high_2014, AuAu_200_production_2016, AuAu200_production2_2016, st_upc, P16id (2014) and P17if (2016) MuDst 
(8) dAu200_production_2016, st_physics, st_fms, P17id, 17133066-17141005, MuDst - st_fms done
(9) production_pp200trans_2015, st_ssdmb, P16id_SL19c
(10) AuAu_200_production_2016, st_mtd, P16ij_SL16j

cuAu_production_2012 P19ic SL19c -
problem found with production. No need to restore now.

New request:  

4. Embedding 
Large datasets:
20192604 LFSUPC strangeness in Run18 AuAu 27 GeV: Lambda, anti-Lambda and Omega flat production done. Producing exponential sample on Cori
20201501 SPIN additional Run12 pp 500 GeV embedding: PA working on the setup
20215301 SPIN Run15 pp200 transverse Production done by PWG, now under review

Small datasets:

20201503 JetCorr pi/K/p in Run18 Isobar 200 GeV
20180201 LFSUPC, He3 in AuAu 39 GeV reproduction: ticket #2561. Improvement seen in >= SL19b
20185201
 
LFSUPC
 rho0, f0, f2, omega, eta Run17 AuAu 54.4 GeV Not started
20192604 LFSUPC strangeness in Run18 AuAu 27 GeV - Under production
20193802 
BulkCorr neutral and charged K* in Run18 AuAu 27 GeV - Test sample produced
20193803 BulkCorr neutral and charged K* in Run14 AuAu 200 GeV
Not started
20201901-04 HF
 Jpsi and Psi(2S) (Pythia8) in Run17 pp510 st_zerobias: on hold until the data is reproduced with correct chain options
20202801, LFSUPC pi/k/p in Run18 AuAu 3.0 GeV (FXT) - second full sample produced under SL19e. nHits mismatch under investigation
20202802, LFSUPC p/d/t/he3 in Run18 AuAu 3.0 GeV (FXT) - second full sample produced under SL19e
20202803, LFSUPC he3 in Run18 AuAu 27 GeV 
Not started
20202804, LFSUPC he3/antihe3 in Run17 AuAu 54 GeV 
Not started
20203001, BulkCorr p/pbar in Run12 pp 200 GeV 
Test sample produced
20203002, LFSUPC Ks/Lambda/AntiLambda in Run14 AuAu200 GeV HFT 
 Under preparation
20203003, LFSUPC pentaquark in Run14 AuAu 200 GeV HFT 
 Under preparation
20203005-09, LFSUPC H3L/H4L/Lambda in Run18 AuAu 3.0 GeV (FXT) 
- Test sample produced
20203010, LFSUPC pi/k/p in Run18 AuAu 27 GeV 
Done
20203701, LFSUPC H3L 3-body decay in Run18 AuAu 3.0 GeV (FXT) Done
20203801-04, LFSUPC H3L/H4L/Lambda in Run18 AuAu 7.2 GeV (FXT) Not started
20210201, LFSUPC K0S in Run18 AuAu 3.0 GeV (FXT)


Know issues: 

(1) 2016-2018 data-embedding disagreement in nHit.
BulkCorr: Run17 AuAu54 proton re-produced with new TpcRs, better in 0-5% but overcorrect in other centrality bins. Overall effect on proton tracking efficiency less than a couple%